# Towards Unifying Interpretability and Control: Evaluation via Intervention

### v1

@@ 3.2 Comparative Evaluation Across Methods and Models
Given the overall lack of standardized evaluation of mechanistic interpretability methods, we intend for this work to serve as a starting point for systematic evaluation by evaluating methods in simple, easy-to-measure contexts. In particular, we think of our evaluations as measuring a kind of upper bound for these methods: in the easiest of settings, how well do existing methods work?
Explanation Correctness
@@
We first propose metrics to evaluate the correctness of explanations and interventions. More specifically, to test whether a single feature of an explanation $z_{i}$ is correct, we intervene on that feature to produce $z_{i}^{\prime}$ and decode $z^{\prime}$ to $\hat{x}^{\prime}$. We can then hypothesize about the corresponding change in the model’s output when generating from $\hat{x}^{\prime}$ beforehand, as it should match the intervention made to produce $z^{\prime}$. For example, if feature $i$ encodes the concept “references to Paris,” increasing the value of $z_{i}$ should result in increases to references of Paris in the model’s output. From this, we propose a metric of Intervention Success Rate, which measures if increasing the activation of feature $i$ by increasing $z_{i}$ results in the appropriate increase of the feature $i$ in the model’s output. To evaluate a continuous relaxation of this, we can also similarly measure the probability assigned to tokens relating to feature $i$. As such, even if the model’s output does not directly reflect interventions made to $z_{i}^{\prime}$ due to sampling, we can measure if increasing the activation of $i$ results in any change to the model’s output at all. We refer to this metric as Intervened Token Probability. Importantly, both of these metrics can be thought of as measuring the causal fidelity of the features highlighted by explanations.
Usefulness of Intervention Methods
[…]
An Open-ended Evaluation Dataset.
@@ 4 Experiments
In this section, we evaluate the five methods on Gemma2-2b Team et al. (2024), Llama2-7b Touvron et al. (2023), and GPT2-small Radford et al. (2019). We first measure the completeness of the explanations produced in Section 4.2. We then benchmark the methods on the metrics described in Section 3.2, providing quantitative results in Sections 4.3 and 4.4 and qualitative results in Section 5. Finally, we present an analysis of the empirical directional similarities between methods in Section 4.5, and analyses of the efficacy of methods at different model depths in Appendix A.6.
@@ 4.1 Implementation Details
Intervention Topics. We pick 10 intervention topics that all relate to references to specific words or phrases: {‘beauty’, ‘chess’, ‘coffee’, ‘dogs’, ‘football’, ‘New York’, ‘pink’, ‘San Francisco’, ‘snow’, ‘yoga’}. These simple, low-level features are ideal for evaluation through intervention for four key reasons: first, measuring the presence of a word or phrase is much easier than measuring a high-level abstract concept such as sycophancy, second, these features were available across many layers for the Gemma2-2b and GPT2-small sparse autoencoders, third, the features necessarily exist in the Logit Lens unembedding dictionary, and finally, datasets that are labelled for the presence of these features are very straightforward to collect for generating steering vectors and probes. As such, we can easily compare intervention on these features across methods and also easily measure intervention success by checking if the given word/phrase exists in the model’s output.
@@
Steering vectors and probing. We implement steering vectors as described in Rimsky et al. (2023) with a few simple modifications. Where in Rimsky et al. (2023), the difference between contrastive pairs is taken only at the last token, we find that averaging across the token dimension and taking the difference between those averages yields much better results. This is due to the fact that in Rimsky et al. (2023), the only difference between representations occurred in the token position of the answer letter, or the last token; however, in our case the information related to the intervention feature could be present at any token. Contrastive data pairs were generated by prompting ChatGPT for 200 pairs sentences or phrases about the desired intervention topic and about other topics within the same genre. For example, we used the prompts “Can you please generate 100 short sentences or phrases with the color ‘pink’ in them?” to create positive samples and “Can you also generate 100 phrases or sentences with other colors (not pink) in them?” to create negative ones. We also use these contrastive pairs to train the linear probes, using the implementation from Chen et al. (2024). All probes reached train and test accuracies of 100% across all models and intervention topics.
Sparse autoencoders. We focus specifically on sparse autoencoders trained to interpret the residual stream of transformer models. We use the SAELens library from Bloom (2024b) for GPT2-small and the Gemma Scope SAEs released by Google in Lieberum et al. (2024) for Gemma2-2b. For both models, SAE feature labels were found via Neuronpedia Lin & Bloom (2023), which allows users to interact with, explore, and search for auto-interpretation labelled learned features. Note that the Gemma SAEs are of the JumpReLU architecture from Rajamanoharan et al. (2024b) whereas the GPT2-small SAEs are standard ReLU + L1 regularization SAEs, as noted in Bloom (2024a).
4.2 Sanity Checking Explanation Reconstructions
We also measure the coherence of the outputs produced by replacing $x$ with $\hat{x}$, as shown in Appendix Figure 7. We find that the coherence of the outputs generated by the reconstructed latents generally matches the coherence of the clean model outputs. We use a deviation of ±$\pm 1$ around the mean of clean output coherence scores as a threshold for future evaluations, shown in the dashed lines.
Figure 2: Evaluation of the Intervention Success Rate for each method. Note that normalized edit distance is a proxy for intervention intensity that is comparable across methods. Logit Lens outperforms all other methods across all models.
4.3 Intervention Success Across Models
As described in Section 3.2, in order to evaluate the correctness of explanations, we measure the causal effects of intervening on specific features of each explanation. For a given feature or intervention topic $i$, we see if increasing the activation of that feature results in an increase of the feature in the model’s output in the top row of Figure 2. In order to compare across methods, which all have different explanation feature spaces and scales, we measure the success of interventions as a function of the norm of the distance between the edited latent representation $\hat{x}^{\prime}=g(Edit(f(x)))$ and the original latent representation $x$: $||\hat{x}^{\prime}-x||/||x||$ as done in Section 4.2. Results for intervened token probability are in Appendix A.4.
We find that across all three models, Logit lens and Tuned lens have the highest intervention success rate, regardless of the normalized edit distance. While increasing intervention strength and editing the latent representations more for steering vectors and probes results in greater intervention success for GPT2-small and LLama2-7b, the same is not true for Gemma2-2b, for which intervening with steering vectors and probes seems to have minimal effect. Furthermore, we find that intervening with sparse autoencoders is moderately successful, but due to the large reconstruction error for GPT2-small, even with intervention, the edit distance is much greater than that needed for Logit lens and Tuned lens. In general, we believe the lower performance of Sparse Autoencoders is due to heavy noise in the labels of features. For example, a feature labelled ‘references to coffee’, is sometimes actually a feature that encodes for references to ‘beans’ and ‘coffee beans’, and thus only sometimes increases mentions of ‘coffee’. Probes and steering vectors also have suboptimal performance, often due to learning of spurious correlations in the training data rather than the true intervention feature.
@@ 4.3 Intervention Success Across Models
When measuring token probability rather than output presence, as explored in Appendix A.4, we see that intervention with all methods across all models increases the probability of intervention-related tokens, even if the intervention does not succeed. We also note that there is a significant difference between the order of magnitude of the intervened token probability for sparse autoencoders, around $10e^{-5}$ and the rest of the methods, which range from $10e^{-4}$ to 0.5.
4.4 Effects of Intervention on Output Quality
We next measure the coherence of the intervened output text produced by each method to ensure that intervention through mechanistic interpretability methods is possible in applications and does not damage the utility of the model. We measure coherence as described in Section 3 as a function of the normalized latent edit distance $||\hat{x}^{\prime}-x||/||x||$ and with respect to the intervention success rate in Figure 4. We visualize the mean of coherence scores for the clean model outputs with solid black horizontal lines, the same as those shown in Figure 7, with a buffer of ±$\pm 1$ around the mean in dashed lines. We also consider a prompting baseline, where we simply prompt the language model to constantly talk about the intervention topic, to better understand the optimal coherence possible while satisfying the intervention. This is shown by the teal stars in 4. Given that only Llama2-7b was instruction-tuned out of the models evaluated, prompting was less successful for Gemma2-2b and infeasible for GPT2-small. Also note that the intervention success rate approaches near 100% with prompting as the number of generated tokens increases; however, seeing as we only generate 30 tokens, the success rate is slightly lower than expected.
We find that for GPT2-small and Llama2-7b, intervention with most methods results in a sharp decline in text coherence, with Logit Lens and Tuned Lens having the most dramatic fall-off. In fact, we find that even with a full point drop in coherence from the baseline models’ coherence (denoted with the black lines), the representations can barely be edited by any methods except for Gemma2-2b. While at first, this seems quite concerning, as this would prevent the real-world application of these methods to scenarios necessitating high model performance, viewing coherence as a function of intervention success rate tells a different story, as shown in Figure 4. In fact, we see that Logit Lens and Tuned Lens have relatively good coherence-intervention tradeoffs for GPT2-small and LLama2-7b, with success rates of around 0.5 and 0.6 respectively when considering outputs within one point of deviation from the mean coherence score of the clean model. All other methods exhibit much less desirable Pareto curves. We also note that as the models increase in size and performance, with GPT2-small being the smallest and worst language model and Llama2-7b being the largest and most performant, the Pareto curves are shifted up and to the right more for most methods, suggesting that targeted intervention is more possible as models become more capable.
Figure 3: Intervened model output coherence measured with respect to normalized edit distance. The solid horizontal line shows the mean of coherence scores for the clean model outputs, and the dashed lines show $\pm$1 standard deviation around the mean.

### v2

@@ 3.2 Evaluation Across Methods and Models
Given the overall lack of standardized evaluation of mechanistic interpretability methods, we intend for this work to serve as a starting point for systematic evaluation by testing methods in simple, easy-to-measure contexts. In particular, we think of our evaluations as measuring a kind of upper bound for these methods: in the easiest of settings, how well do existing methods work?
@@
Explanation Correctness. We first propose metrics to evaluate the correctness of explanations and interventions. More specifically, to test whether a single feature of an explanation $z_{i}$ is correct, we intervene on that feature to produce $z_{i}^{\prime}$ and decode $z^{\prime}$ to $\hat{x}^{\prime}$, which should generate text that matches the intervention made to produce $z^{\prime}$. For example, if feature $i$ encodes the concept “references to Paris,” increasing the value of $z_{i}$ should result in increases to references of Paris in the model’s output. From this, we propose a metric of Intervention Success Rate, which measures if increasing activation $z_{i}$ results in the appropriate increase of the feature $i$ in the model’s output. To evaluate a continuous relaxation of this, we can also similarly measure the probability assigned to tokens relating to feature $i$. As such, even if the model’s output does not directly reflect interventions made to $z_{i}^{\prime}$ due to sampling, we can measure if increasing the activation of $i$ results in any change to the model’s output at all. We refer to this metric as Intervened Token Probability. Importantly, both of these metrics can be thought of as measuring the causal fidelity of the features highlighted by explanations.
Figure 2: Evaluation of the Intervention Success Rate with respect to edit distance for each method on four models for the simple intervention topics. Note that normalized edit distance is a proxy for intervention strength that is comparable across methods. Logit Lens generally outperforms all other methods.
Usefulness of Intervention Methods. While intervention is a useful method for evaluating the correctness of explanations, it is also a desideratum of its own and a frequent motivation for many explanation methods. For example, methods are often developed for the purpose of de-biasing model outputs or increasing model safety, either by localizing bad behavior or identifying it at inference time, thus allowing for targeted edits to be made. However, a lack of this direct connection between interpretation and model intervention has led to illusory results in prior literature (Hase et al., 2024; Wattenberg & Viégas, 2024). By directly and explicitly measuring how effective interpretability methods are at allowing for targeted intervention or steering, we can avoid such failure cases. Importantly, intervention is only useful if the language model retains its overall performance and still satisfies the purpose of the query as well as the intervention. Thus, we want to evaluate whether interpretability methods can steer model outputs towards feature $i$ without damaging the model. We define Coherence as the grammatical correctness, consistency, and relevance to the prompt of the generated text, which can be measured by querying an appropriate oracle, such as a human or strong LLM. Similarly, we can also measure the Perplexity of the intervened outputs with respect to a strong language model. In practice, we use Llama3.1-8b for both of these metrics, as it is reasonable sized, high-performing, and open source, allowing for the measurement of perplexity. We compare coherence scores given by Llama3.1-8b to those generated by human raters as well as a rules-based grammar checker to ensure efficacy of our LLM-as-a-judge setup in Table 1.
@@ 4 Experiments
In this section, we evaluate the four interpretability methods on our metrics from Section 3.2. We also provide case studies of intervention on more complex and safety-relevant concepts in Section 4.4. Finally, we present an analysis of the empirical alignment between methods in Section 4.5 and a comparison of the strengths and weaknesses of each method in Section 4.6. Additional experiments relating to latent reconstruction error and intervention efficacy across model layers are in Appendix A.1 and A.4.
@@ 4.1 Implementation Details
Intervention Topics. We choose 10 intervention topics that all relate to references to specific words or phrases: {‘beauty,’ ‘chess,’ ‘coffee,’ ‘dogs,’ ‘football,’ ‘New York,’ ‘pink,’ ‘San Francisco,’ ‘snow,’ ‘yoga’}, generalizing ‘Golden gate Claude’-style interventions. These simple, low-level features are ideal for evaluation through intervention for four key reasons: first, measuring the presence of a word or phrase is much easier than measuring a high-level abstract concept such as sycophancy, second, these features were present in the pretrained and labelled sparse autoencoders we studied, third, the features necessarily exist in the Logit Lens unembedding dictionary, and finally, datasets that are labelled for the presence of these features are very straightforward to collect for generating steering vectors and probes. As such, we can easily compare interventions on these features across all interpretability methods and measure intervention success by checking if the given word/phrase exists in the model’s output.
Steering vectors and probing. We implement steering vectors with Contrastive Activation Addition (CAA) Rimsky et al. (2023) with a few simple modifications. Where in CAA, the difference between contrastive pairs is taken only at the last token, we find that averaging across the token dimension and taking the difference between those averages yields much better results. This is due to the fact that in CAA, the only difference between representations occurred in the token position of the answer letter, or the last token; however, in our case the information related to the intervention feature could be present at any token. Example contrastive data pairs were hand-generated by the authors and then used to prompt ChatGPT to create a total of 200 pairs of sentences. All data was verified by the authors and is made available in the accompanying codebase. These contrastive pairs were also used to train the linear probes, using the implementation from Chen et al. (2024). All probes reached train and test accuracies of 100% across all models and intervention topics.
Sparse autoencoders and supervised dictionaries. We focus specifically on sparse autoencoders trained to interpret the residual stream of transformer models. We use the SAELens library from Bloom (2024) for GPT2-small and Llama3-8b and the Gemma Scope SAEs Lieberum et al. (2024) for Gemma2-2b. SAE feature labels were found via Neuronpedia Lin & Bloom (2023), which allows users to search through fully trained SAEs and their auto-interpretation labelled features. We also evaluate the Rank-1 Representation Finetuning (ReFT-r1) supervised dictionaries released by Wu et al. (2025), which have features that directly correspond to the SAE features for Gemma2-2b. Note that dictionaries were only released for layer 20 of Gemma2-2b, so we cannot present evaluation for other layers or models.
4.2 Intervention Success Across Models
As described in Section 3.2, in order to evaluate the correctness of explanations, we measure the causal effects of intervening on specific features of each explanation. For a given feature or intervention topic $i$, we see if increasing the activation of that feature results in an increase of the feature in the model’s output for the ten simple intervention topics. In order to compare across methods, which all have different explanation feature spaces and scales, we measure the success of interventions as a function of the norm of the distance between the edited latent representation $\hat{x}^{\prime}=g(Edit(f(x)))$ and the original latent representation $x$: $||\hat{x}^{\prime}-x||/||x||$. Results for intervention success rate are shown in Figure 2 and results for intervened token probability can be found in Appendix A.5.
Across methods and models, we find that by increasing intervention strength, or the magnitude of the edit to the latent representation, intervention success rate first improves and then levels out, as expected. However, we unexpectedly find that Logit lens and Tuned lens generally have the highest intervention success rate, regardless of the normalized edit distance, except when compared to ReFT-r1 on Gemma2-2b. Furthermore, we find that SAEs, probes, and steering vectors require significantly larger edits in order to achieve reasonable intervention success. Note that the minimal edit distance for SAEs is nonzero, as SAE reconstruction incurs a significant error, as explored in Appendix A.1. In general, we believe the lower performance of SAEs is due to heavy noise in the labels of features. For example, a feature labelled ‘references to coffee’, is sometimes actually a feature that encodes for references to ‘beans’ and ‘coffee beans’, and thus only sometimes increases mentions of ‘coffee’. Probes and steering vectors also have suboptimal performance, often due to learning of spurious correlations in the training data rather than the true intervention feature.
4.3 Effects of Intervention on Output Quality
We next measure the coherence of the intervened output text produced by each method to ensure that intervention through interpretability methods is possible without damaging the utility of the model. We measure coherence as described in Section 3 as a function of the intervention success rate in Figure 3 to characterize the tradeoff between intervention success and output coherence. Results for coherence as a function of normalized latent edit distance, $||\hat{x}^{\prime}-x||/||x||$, are in Appendix A.3. We visualize the mean of coherence scores for the clean model outputs with solid black horizontal lines, the same as those shown in Figure 7, with a buffer of ±$\pm 1$ around the mean in dashed lines. We also consider a prompting baseline, where we simply prompt the language model to talk about the intervention topic, to better understand the optimal coherence possible while satisfying the intervention. This is shown by the teal stars in Figure3. Prompting was infeasible for GPT2-small as it was not instruction tuned. Also, note that the intervention success rate approaches 100% with prompting as the number of generated tokens increases; however, seeing as we only generate 30 tokens, the success rate may be lower than expected.
Our experiments reveal that while interpretability methods may seem to provide reasonable tradeoffs between intervention success and coherence at first glance, they all underperform the simplest baseline of just prompting the model. Furthermore, Logit lens and Tuned lens significantly outperform all other methods when intervening on these simple topics, with intervention success rates of around 0.5 and 0.6 respectively for outputs within one point of deviation from the mean coherence score of the clean model. All other methods exhibit much less desirable Pareto curves, regardless of model size or intervention feature.
Table 1: Correlation between human raters (left) and an LLM rater (Llama3-8b) for coherence or a rules-based grammar checker (right). All three raters are highly correlated with one another.
LLM Rater vs
Human Rater
LLM Rater vs
Error Checker
Pearson r $r^{2}$ Pearson r $r^{2}$
Llama3-8b 0.94 0.75 -0.96 0.92
Llama2-7b 0.80 0.68 -0.85 0.73
Gemma2-2b 0.80 0.67 -0.78 0.75
GPT2-sm 0.71 0.67 -0.86 0.74
Verifying Coherence. In order to validate the coherence scores generated through our LLM-as-a-judge setup with Llama3-8b, we verify the coherence scores with human raters. Participants blindly rated 100 outputs for each model, and we measured the correlation between these human ratings and LLM ratings, as shown in Table 1. We find high consistency between both, with particularly high correlation coefficients and $r^{2}$ values for the larger models.
We also check the validity of the coherence ratings by comparing with an alternative metric that measures the number of grammatical errors in the intervened output via a rule-based grammar checker. In particular, we use LanguageTool to determine the number of errors in each output, which has thousands of rules relating to grammar, typos, capitalization errors, and more. As expected, there is a high negative correlation between the two, indicating that outputs with more errors are less coherent. However, we note that the number of grammatical errors is not an ideal metric, as it does not assess whether the text generation pertains to the prompt, which an LLM rater can do.
Qualitative Examples. We present examples of intervention outputs in Figure 5 for the feature ‘yoga,’ with more examples in Appendix A.7. We highlight outputs where intervention succeeded with minimal degradation in coherence in “Optimal Intervention Strength” (left column) as well as generations from the highest intervention strength tested (right column) “Excessive Intervention.” Note that intervention results in repetition at very high intervention strengths for all methods; however, only Logit Lens and Tuned Lens result in repetition of tokens related to ‘yoga.’
Figure 4: Relationship between intervention success rate and coherence for three complex features: religious references (top), gendered language (middle), and French language (bottom) for Gemma2-2b (left) and Llama3-8b (right).
4.4 Intervention on Complex Features
While the aforementioned simple features allow for rigorous evaluation across methods, in practice, users often want to control or steer much more complex concepts. To investigate the feasibility of interventions in more interesting and realistic settings, we present results for three more complex concepts: (1) religious speech, with direct or implicit references to a given religion, (2) gendered-language, or the ability to preferentially generate text related to a specific gender, and (3) the French language, where the model should generate text in French even when given an English prompt. These concepts were chosen from features known to exist in the pretrained and labeled SAEs we evaluated, which had a female gendered-language concept and a Christianity- and Islam-related religion concept. [^2] ReFT-r1 did not have a feature that directly corresponded to French language, so we consider the closest successful feature available: “French connective and referential pronouns.” Intervention with Logit lens was performed by selecting ten relevant tokens as the features for the gender and religion concepts, such as {‘she’, ‘her’, ‘herself’} and {‘Christ’, ‘Allah’, ‘holy’}, but intervention could not be performed for the French concept. Intervention success is measured with keyword detectors for the first two concepts and the NLP-based language detector package polyglot Al-Rfou (2022) for the French concept.
As shown in Figure 4, we find that even for these more complex concepts, prompting generally outperforms interpretability-based interventions, particularly for the French language concept. Furthermore, no other method performs consistently across models or features, highlighting the unreliability of existing interpretability methods for controlling complex and safety-relevant concepts. However, we generally found that interventions made by Logit lens were most successful in incorporating the intervention into a plausible completion of the prompt. For example, one intervened output was “Whenever I’m outdoors and in nature, I always have a camera with me. I love taking pictures of God’s beautiful creation. I’m a big fan of Jesus and I love spending…”, where we can see that the religion concept is integrated well with the response to the prompt. For most other methods, such as SAEs, interventions either immediately degraded model performance by inducing severe repetition (e.g. “I had a conversation recently about the last and final prophet of the last and final of the last and the last and the last and”) or they were either unrelated to the prompt or simply incorrect (e.g. “Tell me an interesting fact about a musical instrument. For example, did you know that the piano is actually a Christian Muslim?”).
4.5 Intervention Similarity Between Methods
Figure 6: Cosine similarity between methods’ intervention directions in model latent space across methods.
Given that these methods all result in linear edits that should correspond to the same feature, ideally their interventions should all point in the same direction in the model’s latent space. We evaluate the empirical similarity between methods by measuring the cosine similarity between edit directions, $\hat{x}^{\prime}-x$, for each intervention topic. The average cosine similarity between these vectors for each intervention topic is shown in Figure 6.
We find that Logit Lens and Tuned Lens are highly similar, as expected. Similarly, steering vectors and probe weights tend to lie in similar directions, likely due to the same underlying data used to train both. Most interestingly, we find that sparse autoencoders tend to intervene in somewhat similar directions to steering vectors and probes and have near orthogonal directions to Logit Lens and Tuned Lens, even when interventions succeed for all methods. We speculate that sparse autoencoders may be more similar to probes and steering vectors because the three methods may have a bias toward representing past information and tokens, due to their training and labelling algorithms, also noted by Gur-Arieh et al. (2025). Logit lens and Tuned lens, on the other hand, are designed to reveal information about the next token specifically, given that they are early-decoding strategies and thus may contain more information about model outputs rather than inputs.
4.6 High-Level Comparison of Methods
We present the following characterization of the strengths and weaknesses of the interpretability and steering methods.
Logit Lens is easy to use, requires no training, and maps features directly to vocabulary tokens, making it highly interpretable, and we find that it generally has high causal fidelity as well. However, its predefined, static features are limited and rudimentary, preventing its utility in many real-world or safety-critical applications. Tuned Lens shares these traits but requires an additional learned affine transformation, although many are already open-sourced. Sparse autoencoders learn a wide variety of both low- and high-level features; however, the post-hoc labeling of features is often highly erroneous and leads to lower causal fidelity in practice. They are also extremely data- and compute-intensive, and there is no guarantee that any given desired feature will exist in the SAE’s dictionary. Supervised dictionaries can contain any arbitrary feature but also require significant data and compute to learn, along with an exhaustive list of relevant features. Importantly, they can be trained with intervention in mind, increasing their utility for steering while maintaining interpretability. Probing and steering vectors also allow for feature specification but are prone to learning spurious correlations, leading to low causal fidelity, and both require labelled data.
As such, we believe that lens-based methods are most useful for providing high-fidelity explanations but are likely not effective solutions for steering in real applications, for which steering vectors and prompting are more promising but require careful oversight and refinement to ensure efficacy due to their uninterpretability. Finally, supervised dictionaries seem to hold a nice balance between interpretation- and control-based methods if users are not compute- and data-bound, proving to be more promising that unsupervised methods such as SAEs.
5 Conclusion
While interpretability methods show great promise in understanding large language models, the correctness of their explanations is less clear. Do these explanations reveal truth about model computation or simply fool human researchers? We believe that systematic benchmarking of explanations is critical to answer this question. Our work makes progress towards this goal, and answers this question somewhat negatively, showing that current explanations are less accurate than expected. Our work also raises questions regarding the utility of such methods, as we find that prompting outperforms current interpretability methods in its ability to steer models, without requiring any training, data, or access to model weights. We hope future work can address these shortcomings of current methods, paving way toward interpretability methods that are faithful and provide actionable insights for improving and controlling models.
Acknowledgements
Impact Statement
This paper aims to improve the evaluation of interpretability methods, ensuring that they are both faithful and useful in practice. We are particularly interested in facilitating discussions around how to make interpretability methods more reliable to improve model fairness and safety, both of which can have significant societal implications. Ultimately, we hope that clear evaluation of interpretability methods will ensure that methods produce desired results, decreasing unintended consequences due to spurious correlations or unfaithful explanations. However, we note that interpretability can just as easily be used to make models more harmful than helpful.
Appendix A Appendix
A.1 Additional Evaluations: Sanity Checking Explanation Reconstructions
@@
A.2 Additional Evaluations: Coherence of Method Outputs without Intervention
A.3 Additional Evaluation: Coherence of Intervention with respect to Edit Distance
We measure the coherence of the intervened output text produced by each method to ensure that intervention through interpretability methods is possible without damaging the utility of the model. We measure coherence as described in Section 3 as a function of normalized latent edit distance, $||\hat{x}^{\prime}-x||/||x||$ in Figure 8. We find that even the smallest interventions made with Logit lens and Tuned lens result in significant degradation of model outputs, with a less noticeable dropoff for the other methods. We also plot coherence as a function of the intervention success rate in Figure 3 to characterize the tradeoff between intervention success and output coherence.
Figure 8: Analysis of coherence of the intervened outputs, measured with Llama3.1-8b, as a measure of the edit distance or magnitude of intervention made. Lens-based methods suffer drastic drops in coherence with only small edits.
A.4 Additional Evaluations: Intervention Efficacy Across Model Depth
Figure 9: Analysis of intervention pass rate (left), coherence (middle) and edit distance (right) across all layers of GPT2-sm. We find that intervening at later layers is significantly more effective for Logit and Tuned Lens than earlier interventions, but probes, steering vectors, and SAEs are relatively invariant to the choice of layer.
A.5 Additional Metrics: Intervened Token Probability
Please see Section 3.2 for more details. We measure the probability assigned to tokens relating to feature $i$ when intervening on feature $i$. As such, even if a model’s output does not directly reflect interventions made to $z_{i}^{\prime}$ due to sampling, we can measure if increasing the activation of feature $i$ results in any change to the model’s output at all. We refer to this metric as Intervened Token Probability.
@@ A.5 Additional Metrics: Intervened Token Probability
Results for Intervened Token Probability are shown in Figure 10, where we see that intervention with all methods across all models increases the probability of intervention-related tokens, even if the intervention does not succeed. We also note that there is a significant difference between the order of magnitude of the intervened token probability for sparse autoencoders, around $10e^{-5}$ and the rest of the methods, which range from $10e^{-4}$ to 0.5.
Figure 10: Evaluation of intervention success with respect to the probabilities of the tokens corresponding to the features intervened on for each method. Note that normalized edit distance is a proxy for intervention intensity that is comparable across methods.
A.6 Additional Metrics: Perplexity
As described in Section 3.2, we evaluate the perplexity of the intervened generated text to measure the utility of interpretability methods for targeted intervention in 11. We measure this perplexity with respect to a stronger language model than the one studied, in this case with Llama3.1-8b.
Figure 11: Analysis of perplexity of the intervened outputs, measured with Llama3.1-8b, as an alternative metric to Coherence. We find that perplexity does not align with Coherence, as highly repetitive sequences may have low perplexity despite being incoherent answers to prompts.
A.7 Additional Example Outputs